Three point charges of 3µC, 8µC and 5µC are located at the corners of an equilateral triangle of side 0.4m, calculate the net force on the 8µC?
1
Expert's answer
2019-04-18T09:58:24-0400


The components of a net force on a second charge


"F=\sqrt{F_{12}^2+F_{32}^2-2F_{12}F_{32}\cos 120^{\circ}}"

"=\frac{k q_2}{a^2}\sqrt{q_1^2+q_3^2+q_1q_3}"

"=\frac{9\times 10^9\times 8\times 10^{-6}}{(0.4)^2}\sqrt{3^2+5^2+3\times 5}\times 10^{-6}"

"=3.15\:\rm{N}"